OpenStudy (anonymous):

Convergent/Divergent integral--- Integrate from [negative infinity, 0] (x^2)*(e^x) I'll actually try to quickly post below how far I've gotten, to see if I'm right so far....

OpenStudy (anonymous):

\[\lim_{t \rightarrow -\infty} \int\limits_{t}^{0} x^2e^xdx\]

OpenStudy (anonymous):

\[\lim_{t \rightarrow -\infty} -1/2 - t^2e^t + 1/2(e^t)\]

OpenStudy (turingtest):

I don't think you integrated right

OpenStudy (anonymous):

I integrated by parts...

OpenStudy (anonymous):

which gave me = \[x^2e^x - (1/2)\int\limits_{t}^{0}e^xdx\] Everything evaluated at t and 0...

OpenStudy (turingtest):

\[\int x^2e^xdx\]\[u=x^2\implies du=2xdx\]\[dv=e^xdx\implies v=e^x\]\[\int x^2e^xdx=x^2e^x-2\int xe^xdx\]

OpenStudy (anonymous):

ahhh.... maybe I did integrate wrong....hmm. Messed up on the 2xdx part...

OpenStudy (turingtest):

yeah you did some non-existent u-substitution thing in your head I guess..

OpenStudy (turingtest):

so what should the integral be?

OpenStudy (anonymous):

Mkay... so I'm left with \[\lim_{t \rightarrow -\infty} 2 - t^2e^t + 2e^t t - 2e^t\]

OpenStudy (anonymous):

Which leaves me to .... have to use the LH rule? I'm thinking...

OpenStudy (turingtest):

yeah, but you gotta do a little substitution befor you can use it

OpenStudy (turingtest):

\[\lim_{t \rightarrow -\infty} 2 - t^2e^t + 2e^t t - 2e^t\]\[=2-\lim_{t \rightarrow -\infty} t^2e^t + 2e^t t - 2e^t\]let \(n=-t\) and what do you get?